The length of a rectangular garden is 7 feet longer than its width. If the garden's perimeter is 198 feet, what is the area of the garden in square feet?
Answer by addingup(3677)

Length = Width + 7
2L + 2W = 198
substitute for L:
2(W+7) + 2W = 198
2W + 14 + 2W = 198
4W = 184
W = 46 and L = W + 7 = 46 + 7 = 53
Check:
2L + 2W = 198
2(53) + 2(46) =106+92 = 198 Correct